# Onsager-Machlup Functional and Large Deviation Principle for Stochastic Hamiltonian Systems

### v1

@@ 3. Onsager-Machlup functional for stochastic Hamiltonian systems
(3.15) $\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{13}\right\}\big{|}\left\|W^{\sigma}\right\|<\epsilon\right)=1$
for all $c\in\mathbb{R}$, as $\epsilon\to 0$ and $\delta\to 0$.
From the [46], for a general stochastic differential equation:
$\mathrm{d}X_{t}=f(t,X_{t})\mathrm{d}t+g(t)\mathrm{d}W_{t},$
the Onsager-Machlup functional, which accounts for path deviations caused by drift and disturbances, is given by:
@@
$\int_{0}^{1}OM(\varphi,\dot{\varphi})\,{\rm d}t=\int_{0}^{1}\left|{g(t)}^{-1}\cdot\left({\dot{\varphi}_{t}-f(t,\varphi_{t})}\right)\right|^{2}\,{\rm d}t+\int_{0}^{1}{{\rm div}^{g}_{x}f(t,\varphi_{t})}\,{\rm d}t,$
where $\dot{\varphi}:=\frac{\mathrm{d}\varphi(t)}{\mathrm{d}t}$, and ${\rm div}^{g}_{x}f(t,\varphi_{t}):=\mathrm{Tr}.\left({g(t)^{-1}}{\nabla f(t,\varphi_{t})}g(t)\right)$, representes a correction term.
[…]
$\quad OM(\varphi_{q},\varphi_{p})$
$=\int_{0}^{1}\left|\sigma_{q}^{-1}(t)\left(\dot{\varphi}_{q}-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right)\right|^{2}\,{\rm d}t+\int_{0}^{1}\left|\sigma_{p}^{-1}(t)\left(\dot{\varphi}_{p}+\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right)\right|^{2}\,{\rm d}t.$
4. The most probable path in stochastic Hamiltonian systems
The minimum value of the Onsager-Machlup functional $OM(\varphi_{q},\varphi_{p})$ corresponds to the most probable continuous paths $\hat{\varphi}(t):=(\hat{\varphi}_{q},\hat{\varphi}_{p})$ in stochastic Hamiltonian systems. Due to the unique form of the Onsager-Machlup function in such systems:
$\quad OM(\varphi_{q},\varphi_{p})$
$=\int_{0}^{1}\left|\sigma_{q}^{-1}(t)\left(\dot{\varphi}_{q}-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right)\right|^{2}\,{\rm d}t+\int_{0}^{1}\left|\sigma_{p}^{-1}(t)\left(\dot{\varphi}_{p}+\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right)\right|^{2}\,{\rm d}t,$
we can define the Lagrangian $L$ using the Euler-Lagrange equations:
$L=\left|\sigma_{q}^{-1}(t)\left(\dot{\varphi}_{q}-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right)\right|^{2}+\left|\sigma_{p}^{-1}(t)\left(\dot{\varphi}_{p}+\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right)\right|^{2}.$
Through the Hamiltonian principle, also known as the principle of least action, we can directly determine the most probable continuous path $\hat{\varphi}(t)$ that satisfies the following equations:
(4.1) $\begin{cases}d\varphi_{q}(t)=\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q}(t),\varphi_{p}(t))\,{\rm d}t,\\ d\varphi_{p}(t)=-\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q}(t),\varphi_{p}(t))\,{\rm d}t.\end{cases}$
This indicates that despite the presence of random disturbances, the system is most likely to evolve along the classical Hamiltonian trajectory.
Therefore, although stochastic perturbations introduce complexity and uncertainty into Hamiltonian systems, by minimizing the Onsager-Machlup functional, we can still reveal the dominant factors of system behavior, namely the evolution along the most probable path. This discovery is of significant importance in both theoretical research and practical applications. To further validate this conclusion, we will illustrate it through a specific example in the following section.
Example 4.1.
Consider a one-dimensional harmonic oscillator with the Hamiltonian $H(q,p)=\frac{p^{2}}{2m}+\frac{1}{2}kq^{2}$, where $\frac{p^{2}}{2m}$ represents the kinetic energy term, and $V(q)=\frac{1}{2}kq^{2}$ is the potential energy term. Here, $m=1$ denotes the mass, $k=1$ denotes the spring constant, and $q$ and $p$ respectively represent position and momentum. The classical Hamiltonian equations are given by:
[…]
In Fig.1, we observe a clear contrast between the phase space trajectories of the deterministic harmonic oscillator and the stochastically perturbed oscillator. In the absence of perturbations, the deterministic system follows a stable, closed circular orbit, where momentum $p$ and position $q$ evolve periodically, reflecting energy conservation. This circular trajectory represents the steady exchange between kinetic and potential energy, maintaining the Hamiltonian.
Figure 1. Comparison of phase space trajectories between the deterministic (bold solid line) and stochastically perturbed (thin dashed lines) harmonic oscillators. The system is modeled with the following parameters: mass $m=1$, spring constant $k=1$, time step $dt=0.0001$, and total simulation time $T=300$. The deterministic trajectory follows a stable circular orbit in phase space, indicating energy conservation and periodic motion. In contrast, the stochastic system introduces white noise perturbations using the Euler-Maruyama method, leading to diffusive and irregular trajectories that deviate from the original orbit over time.
When white noise perturbations are introduced, however, the trajectory deviates from this ideal path, displaying diffusion and irregularity. As shown in Fig.1, the stochastic oscillator’s trajectory gradually drifts away from the stable orbit, with random fluctuations disrupting the original pattern. This diffusion occurs because the noise introduces energy fluctuations, causing deviations in $p$ and $q$ that prevent strict adherence to the classical oscillator’s path. While some periodic behavior remains, the perturbations induce instability, leading to a trajectory that expands outward in phase space over time.
Nevertheless, we also observe that the stochastic oscillator’s phase space trajectories remain densely concentrated near the deterministic oscillator’s stable circular orbit. This aligns with our theoretical conclusions: the most probable path of the stochastically perturbed system stays close to the stable deterministic trajectory. This reinforces the understanding that, despite random perturbations, the system tends to evolve near the deterministic solution.
Fig.2 provides further insight by displaying the probability distribution of the Hamiltonian $H(q,p)$ for the stochastically perturbed oscillator. Notably, the Hamiltonian distribution $H_{S}$ peaks near $H_{S}=1249.6399$, closely matching the deterministic Hamiltonian value $H_{D}=1250$. This behavior supports our theoretical conclusion that the most probable Hamiltonian value for the stochastically perturbed system coincides with that of the deterministic system. For most of the time, the system’s Hamiltonian remains concentrated near the deterministic value, with large energy fluctuations being rare. This indicates that, despite random perturbations, the system’s overall evolution largely adheres to the conservation properties of the Hamiltonian system, with only local deviations due to noise.
Figure 2. The distribution curve of the Hamiltonian $H(q,p)=\frac{p^{2}}{2m}+\frac{1}{2}kq^{2}$. The system is modeled with the following parameters: mass $m=1$, spring constant $k=1$, time step $dt=0.001$, total simulation time $T=300$, and 5000 simulation runs. The maximum value of the Hamiltonian $H$ is observed at 1249.6399, with a corresponding density of 0.0006. This result illustrates the distribution of the Hamiltonian in the presence of stochastic perturbations, showing that the system’s energy tends to concentrate around specific values in most cases.
In summary, Fig.1 and Fig.2 together illustrate the dynamics of both the classical Hamiltonian system and the stochastically perturbed system, revealing that the effects of random noise on the system’s trajectory and Hamiltonian are limited. Despite the introduction of stochastic perturbations, the system’s behavior remains closely aligned with that of the classical Hamiltonian system, supporting the theoretical conclusions derived from the Onsager-Machlup functional analysis. Next, we apply our results to a high-dimensional example.
Example 4.2.
Consider a simplified three-body system consisting of the Sun, Earth, and Moon. Given that the mass of the Sun ($m_{1}$) is significantly larger than those of the Earth ($m_{2}$) and Moon ($m_{3}$), we approximate the Sun as stationary, with the Earth and Moon moving within the Sun’s gravitational field. The specific parameters and Hamiltonian system are described below:
Sun:
  • Mass: $m_{1}=1.989\times 10^{30}$.
  • Position: $\vec{r}_{1}=(0,0,0)$.
Earth:
  • Mass: $m_{2}=5.972\times 10^{24}$.
  • Position: $\vec{r}_{2}=(x_{2},y_{2},z_{2})$.
  • Initial Position at $t=0$: $(1.496\times 10^{11},0,0)$.
  • Momentum: $\vec{p}_{2}=m_{2}\vec{\dot{r}}_{2}=(p_{x_{2}},p_{y_{2}},p_{z_{2}})$.
Moon:
  • Mass: $m_{3}=7.348\times 10^{22}$.
  • Position: $\vec{r}_{3}=(x_{3},y_{3},z_{3})$.
  • Initial Position at $t=0$: $(1.496\times 10^{11}+3.844\times 10^{8},0,0)$.
  • Momentum: $\vec{p}_{3}=m_{3}\vec{\dot{r}}_{3}=(p_{x_{3}},p_{y_{3}},p_{z_{3}})$.
The Gravitational Constant is given by $G=6.67430\times 10^{-11}$. The Kinetic Energy of the system can be expressed as:
$T:=\frac{1}{2m_{2}}(p_{x_{2}}^{2}+p_{y_{2}}^{2}+p_{z_{2}}^{2})+\frac{1}{2m_{3}}(p_{x_{3}}^{2}+p_{y_{3}}^{2}+p_{z_{3}}^{2}).$
Similarly, the Potential Energy is given by:
$V:=-G\left(\frac{m_{1}m_{2}}{|\vec{r}_{2}|}+\frac{m_{1}m_{3}}{|\vec{r}_{3}|}+\frac{m_{2}m_{3}}{|\vec{r}_{3}-\vec{r}_{2}|}\right),$
where the magnitudes of the position vectors are defined as:
$|\vec{r}_{2}|=\sqrt{x_{2}^{2}+y_{2}^{2}+z_{2}^{2}},\quad|\vec{r}_{3}|=\sqrt{x_{3}^{2}+y_{3}^{2}+z_{3}^{2}},$
$|\vec{r}_{3}-\vec{r}_{2}|=\sqrt{(x_{3}-x_{2})^{2}+(y_{3}-y_{2})^{2}+(z_{3}-z_{2})^{2}}.$
Therefore, the total energy $H$ of the system is:
$H=\frac{1}{2m_{2}}(p_{x2}^{2}+p_{y2}^{2}+p_{z2}^{2})+\frac{1}{2m_{3}}(p_{x3}^{2}+p_{y3}^{2}+p_{z3}^{2})-G\left(\frac{m_{1}m_{2}}{|\vec{r}_{2}|}+\frac{m_{1}m_{3}}{|\vec{r}_{3}|}+\frac{m_{2}m_{3}}{|\vec{r}_{3}-\vec{r}_{2}|}\right).$
In summary, the equations governing the motion of the Earth can be expressed as:
(4.4) $\begin{cases}\dot{x}_{2}=\frac{p_{x_{2}}}{m_{2}},\\ \dot{p}_{x_{2}}=-Gm_{1}\frac{x_{2}}{|\vec{r}_{2}|^{3}}-Gm_{3}\frac{x_{2}-x_{3}}{|\vec{r}_{3}-\vec{r}_{2}|^{3}},\\ \dot{y}_{2}=\frac{p_{y_{2}}}{m_{2}},\\ \dot{p}_{y_{2}}=-Gm_{1}\frac{y_{2}}{|\vec{r}_{2}|^{3}}-Gm_{3}\frac{y_{2}-y_{3}}{|\vec{r}_{3}-\vec{r}_{2}|^{3}},\\ \dot{z}_{2}=\frac{p_{z_{2}}}{m_{2}},\\ \dot{p}_{z_{2}}=-Gm_{1}\frac{z_{2}}{|\vec{r}_{2}|^{3}}-Gm_{3}\frac{z_{2}-z_{3}}{|\vec{r}_{3}-\vec{r}_{2}|^{3}}.\end{cases}$
Similarly, the equations governing the motion of the Moon can be expressed as:
(4.5) $\begin{cases}\dot{x}_{3}=\frac{p_{x_{3}}}{m_{3}},\\ \dot{p}_{x_{3}}=-Gm_{1}\frac{x_{3}}{|\vec{r}_{3}|^{3}}-Gm_{2}\frac{x_{3}-x_{2}}{|\vec{r}_{3}-\vec{r}_{2}|^{3}},\\ \dot{y}_{3}=\frac{p_{y_{3}}}{m_{3}},\\ \dot{p}_{y_{3}}=-Gm_{1}\frac{y_{3}}{|\vec{r}_{3}|^{3}}-Gm_{2}\frac{y_{3}-y_{2}}{|\vec{r}_{3}-\vec{r}_{2}|^{3}},\\ \dot{z}_{3}=\frac{p_{z_{3}}}{m_{3}},\\ \dot{p}_{z_{3}}=-Gm_{1}\frac{z_{3}}{|\vec{r}_{3}|^{3}}-Gm_{2}\frac{z_{3}-z_{2}}{|\vec{r}_{3}-\vec{r}_{2}|^{3}}.\end{cases}$
The normal three-body problem, when influenced by the perturbations from some other distant celestial bodies, can be abstractly modeled as the impact of random disturbances. We may define the three-body problem under random perturbations in the following form:
The Stochastic Equation for Earth’s Motion
(4.6) $\begin{cases}{\rm d}x_{2}=\frac{p_{x_{2}}}{m_{2}}\,{\rm d}t+c_{21}(t)\,{\rm d}W_{21}(t),\\ {\rm d}p_{x_{2}}=-G\left(m_{1}\frac{x_{2}}{|\vec{r}_{2}|^{3}}+m_{3}\frac{x_{2}-x_{3}}{|\vec{r}_{3}-\vec{r}_{2}|^{3}}\right)\,{\rm d}t+c_{22}(t)\,{\rm d}W_{22}(t),\\ {\rm d}y_{2}=\frac{p_{y_{2}}}{m_{2}}\,{\rm d}t+c_{23}(t)\,{\rm d}W_{23}(t),\\ {\rm d}p_{y_{2}}=-G\left(m_{1}\frac{y_{2}}{|\vec{r}_{2}|^{3}}+m_{3}\frac{y_{2}-y_{3}}{|\vec{r}_{3}-\vec{r}_{2}|^{3}}\right)\,{\rm d}t+c_{24}(t)\,{\rm d}W_{24}(t),\\ {\rm d}z_{2}=\frac{p_{z_{2}}}{m_{2}}\,{\rm d}t+c_{25}(t)\,{\rm d}W_{25}(t),\\ {\rm d}p_{z_{2}}=-G\left(m_{1}\frac{z_{2}}{|\vec{r}_{2}|^{3}}+m_{3}\frac{z_{2}-z_{3}}{|\vec{r}_{3}-\vec{r}_{2}|^{3}}\right)\,{\rm d}t+c_{26}(t)\,{\rm d}W_{26}(t).\end{cases}$
The Stochastic Equation for Moon’s Motion
(4.7) $\begin{cases}{\rm d}x_{3}=\frac{p_{x_{3}}}{m_{3}}\,{\rm d}t+c_{31}(t)\,{\rm d}W_{31}(t),\\ {\rm d}p_{x_{3}}=-G\left(m_{1}\frac{x_{3}}{|\vec{r}_{3}|^{3}}+m_{2}\frac{x_{3}-x_{2}}{|\vec{r}_{3}-\vec{r}_{2}|^{3}}\right)\,{\rm d}t+c_{32}(t)\,{\rm d}W_{32}(t),\\ {\rm d}y_{3}=\frac{p_{y_{3}}}{m_{3}}\,{\rm d}t+c_{33}(t)\,{\rm d}W_{33}(t),\\ {\rm d}p_{y_{3}}=-G\left(m_{1}\frac{y_{3}}{|\vec{r}_{3}|^{3}}+m_{2}\frac{y_{3}-y_{2}}{|\vec{r}_{3}-\vec{r}_{2}|^{3}}\right)\,{\rm d}t+c_{34}(t)\,{\rm d}W_{34}(t),\\ {\rm d}z_{3}=\frac{p_{z_{3}}}{m_{3}}\,{\rm d}t+c_{35}(t)\,{\rm d}W_{35}(t),\\ {\rm d}p_{z_{3}}=-G\left(m_{1}\frac{z_{3}}{|\vec{r}_{3}|^{3}}+m_{2}\frac{z_{3}-z_{2}}{|\vec{r}_{3}-\vec{r}_{2}|^{3}}\right)\,{\rm d}t+c_{36}(t)\,{\rm d}W_{36}(t).\end{cases}$
By substituting specific parameters, we compute the value of the Hamiltonian at the initial moment as $H_{0}\approx-2.706\times 10^{33}$. Furthermore, we obtain that the magnitudes of the positional derivatives $\dot{x}_{2}$, $\dot{y}_{2}$, and $\dot{z}_{2}$ for Earth are on the order of $10^{4}$, while the magnitudes of the momentum derivatives $\dot{p}_{x_{2}}$, $\dot{p}_{y_{2}}$, and $\dot{p}_{z_{2}}$ are on the order of $10^{22}$. For the Moon, the magnitudes of the positional derivatives $\dot{x}_{3}$, $\dot{y}_{3}$, and $\dot{z}_{3}$ are on the order of $10^{3}$, and those of the momentum derivatives $\dot{p}_{x_{3}}$, $\dot{p}_{y_{3}}$, and $\dot{p}_{z_{3}}$ are on the order of $10^{20}$.
From a physical perspective, the overall dynamical influence of external factors on the three-body system is relatively minor. Consequently, we set the intensity of random diffusion to be on the order of magnitude of $10^{-5}$ relative to its corresponding quantity. However, during the numerical simulation, to capture the long-time scale of the three-body problem (approximately $10^{18}$ seconds), we performed a time-scale transformation. Specifically, we shorten the time unit by a factor of $10^{16}$ (e.g., 1 second in the simulation now corresponds to $10^{16}$ seconds in reality). Thus, in the numerical simulation, $t\in[0,100]$ is equivalent to $t\in[0,10^{18}]$ in the real world. Since Brownian motion follows $W_{t}\;\sim\;\sqrt{t}$, the coefficient of the noise term needs to be scaled up by a factor of $10^{8}$. In summary, we define $c_{21}(t)=c_{23}(t)=c_{25}(t)=10^{7}$, $c_{22}(t)=c_{24}(t)=c_{26}(t)=10^{24}$, $c_{31}(t)=c_{33}(t)=c_{35}(t)=10^{6}$, and $c_{32}(t)=c_{34}(t)=c_{36}(t)=10^{23}$.
[…]
Figure 3. This represents the distribution curve of the Hamiltonian $H(q,p)$ for the three-body problem. Simulations indicate that the maximum value of the Hamiltonian $H$ is $-2.7059\times 10^{33}$. In the real physical world, the intensity of random perturbations is even smaller. Our numerical simulation results are consistent with the long-term evolution of a three-body system consisting of the Sun, Earth, and Moon, which maintains approximate energy conservation.
5. Large Deviation Principle and Rate Function for Stochastic Hamiltonian Systems
Since our objective is to study the behavior of the trajectories in the stochastic Hamiltonian system as the noise term approaches zero, we simplify Equation (1.1) into the following form for convenience:
(5.1) $\begin{cases}\mathrm{d}q(t)=\frac{\partial H}{\partial p}(q,p)\,\mathrm{d}t+\epsilon\sigma_{q}(t)\,\mathrm{d}W_{q}(t),\\ \mathrm{d}p(t)=-\frac{\partial H}{\partial q}(q,p)\,\mathrm{d}t+\epsilon\sigma_{p}(t)\,\mathrm{d}W_{p}(t),\end{cases}$
where $q(t)$ and $p(t)$ represent the generalized position and momentum variables, respectively, and $H(q,p)$ is the Hamiltonian function describing the system’s energy, typically composed of kinetic and potential energy terms. $W_{q}(t)$ and $W_{p}(t)$ are independent Wiener processes, and the matrices $\sigma_{q}(t)$ and $\sigma_{p}(t)$ denote the diffusion coefficients for the stochastic components. The parameter $\epsilon$ represents the noise intensity. Our primary focus is on analyzing the statistical behavior of this system as $\epsilon\to 0$. Similar to Section 3, our result holds for any finite interval $t\in[0,T]$. For simplicity of notation, we will present the following theorem on the interval $[0,1]$.
@@ 5. Large Deviation Principle and Rate Function for Stochastic Hamiltonian Systems
Assuming that Conditions $C1$ and $C2$ hold, let $X^{\epsilon}:=(q(t),p(t))$ be the solution to the following stochastic Hamiltonian system (5.1). As $\epsilon\to 0$, the most probable path $\hat{\varphi}(t)=(\hat{\varphi}_{q}(t),\hat{\varphi}_{p}(t))$ is given by the deterministic Hamiltonian equations (4.1). For any path $X(t)=(q(t),p(t))$, the probability that the system deviates from the most probable path satisfies the large deviation principle:
(5.2) $\epsilon^{2}\ln\mathbb{P}(X(t)\in A)\approx-\inf_{\varphi\in A}J(\varphi),$
where $\varphi\in A$ denotes an arbitrarily continuous function, and $A\subset\mathbb{R}^{d}$ denote an arbitrary measurable set, and suppose that $\hat{\varphi}(t)\notin A$. Furthermore, the rate function $J(\varphi)$ is given by:
(5.3) $J(\varphi)=\begin{cases}\frac{1}{2}\left(\int_{0}^{1}\left\|\sigma_{q}^{-1}(t)\left(\dot{\varphi}_{q}-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right)\right\|^{2}\,{\rm d}t\right.\\ \left.+\int_{0}^{1}\left\|\sigma_{p}^{-1}(t)\left(\dot{\varphi}_{p}+\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right)\right\|^{2}\,{\rm d}t\right),&\text{if }\varphi-x_{0}\in\mathbb{H}^{1};\\ +\infty,&\text{otherwise}\end{cases}$
with $\sigma_{q}^{-1}(t)$ and $\sigma_{p}^{-1}(t)$ being the inverses of the diffusion matrices $\sigma_{q}(t)$ and $\sigma_{p}(t)$, respectively.

### v2

@@ 4 Most Probable Path
$\dot{\varphi}(t)=J\nabla H(\varphi_{t})-\frac{1}{2}\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}.$ (20)
Proof.
Based on the Ansatz method, we first consider a simplified setting in which the term $\left(D_{y}\sigma(t,\varphi_{t})\right)$ is assumed to depend solely on time $t$ and is denoted by $\mathcal{D}^{\sigma}(t)$. Under this simplification, the Onsager-Machlup functional can be reduced to the following form:
$OM(\varphi,\dot{\varphi})=\int_{0}^{1}{\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}}\,{\rm d}s-\int_{0}^{1}\mathrm{Tr}\left(\mathcal{D}^{\sigma}(s)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s.$ (21)
To facilitate the optimization, we introduce the auxiliary variable
$A(t):=\sigma^{-1}(t,\varphi_{t})\left(J\nabla H(\varphi_{t})-\dot{\varphi_{t}}\right),$
which reformulates the Onsager-Machlup functional as
$OM(A)=\int_{0}^{1}\left(\|A(s)\|^{2}-\mathrm{Tr}(\mathcal{D}^{\sigma}(s)A(s))\right)ds.$
Since $\mathcal{D}^{\sigma}(t)$ is independent of $\varphi$, this expression corresponds to a variational problem where the integrand depends quadratically and linearly on the auxiliary variable $A(t)$. Given that the Frobenius norm is both differentiable and convex, we can compute the extremum by taking the first variation with respect to $A$.
For each fixed $t$, we define the Lagrangian function
$\mathcal{L}(A):=\|A\|^{2}-\mathrm{Tr}(\mathcal{D}^{\sigma}A),$
where $\mathcal{D}^{\sigma}(t):=\mathcal{D}^{\sigma}(t)_{ijk}$ denotes a time-dependent rank-3 tensor and $A:=A_{j}$ is an $n$-dimensional vector. The first variation of $\mathcal{L}$ in the direction of $\delta A$ is given by
$\delta\mathcal{L}$ $=2A^{\top}\delta A-\mathrm{Tr}(\left(\mathcal{D}^{\sigma}\right)^{\top}\cdot\delta A)$
$=2\sum_{j=1}^{n}A_{j}\delta A_{j}-\sum_{i,j}\mathcal{D}^{\sigma}_{jii}\delta A_{j}$
$=\sum_{j=1}^{n}\left(2A_{j}-\sum_{i}\mathcal{D}^{\sigma}_{jii}\right)\delta A_{j},$
which implies that the unique minimizer is
$A^{*}_{j}=\frac{1}{2}\sum_{i}D_{jii}.$
Substituting back the definition of $A$, we obtain the expressions for the most probable path $\varphi(t)$ in the form of a drift-corrected Hamiltonian system
@@
$\dot{\varphi}(t)=J\nabla H(\varphi_{t})-\dfrac{1}{2}\sigma(t,\varphi_{t})\sum_{i}\left({\mathcal{D}^{\sigma}}(t)\right)_{\cdot ii}.$
Based on the Ansatz method, we now assume that the spatial derivatives of the diffusion coefficients, $D_{y}\sigma(t,\varphi_{t})$ is equal to their evaluations at a reference path $\psi(t)=(\psi_{\theta}(t),\psi_{I}(t))$, namely,
$D_{y}\sigma(t,\varphi_{t})=D_{y}\sigma(t,\psi_{t}).$
We further assume that the reference path $\psi(t)$ satisfies the following drift-corrected Hamiltonian system:
$\dot{\psi}(t)=J\nabla H(\psi_{t})-\frac{1}{2}\sigma(t,\psi_{t})\sum_{i}\left({D_{y}}\sigma(t,\psi_{t})\right)_{\cdot ii}.$
Under this assumption, we may again abbreviate the tensor $D_{y}\sigma(t,\psi_{t})$ by the time-dependent third-order tensors $\mathcal{D}^{\sigma}(t)$, which is treated as independent of the variable $\varphi=\left(\varphi_{I},\varphi_{\theta}\right)$.
Following the same steps as in the earlier derivation, we obtain that the most probable path $\varphi(t)$, which minimizes the Onsager-Machlup functional, satisfies the following system:
$\dot{\varphi}(t)=J\nabla H(\varphi_{t})-\frac{1}{2}\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\psi_{t})\right)_{\cdot ii}.$
Since both $\varphi(t)$ and $\psi(t)$ satisfy the same system with identical initial conditions, the uniqueness of solutions to ordinary differential equations implies that $\psi(t)=\varphi(t)$. Therefore, the path $\varphi(t)$ that minimizes the Onsager-Machlup functional - the most probable path - must satisfy the following drift-corrected Hamiltonian system:
$\dot{\varphi}(t)=J\nabla H(\varphi_{t})-\frac{1}{2}\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}.$
This system represents a deterministic perturbation of the classical Hamiltonian equations, where the correction terms arise from the effect of multiplicative noise. In particular, the additional term involving $\frac{1}{2}\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}$ quantify the influence of the multiplicative noise geometry (represented through the spatial derivative tensors of the diffusion matrices) on the most probable evolution path. Under certain regularity assumptions, such as the invertibility of $\sigma$, as well as the smoothness of the Hamiltonian $H\in C_{b}^{3}$, the solution corresponding to the most probable path $\varphi(t)$ is uniquely determined by the initial conditions and the modified dynamics.
Moreover, we observe that the most probable path of the stochastic Hamiltonian system (20) does not, in general, preserve the Hamiltonian structure. To address this issue, we introduce Condition $(C3)$, which is a slightly stronger version of Condition $(C2)$. In Condition $(C3)$, the symmetric matrix assumed in $(C2)$ is further restricted to be diagonal. This modification is both natural and intuitive: while symmetry is a necessary requirement in the derivation of the Onsager-Machlup functional, the preservation of Hamiltonian structure inherently demands antisymmetry. Furthermore, to ensure the conservation of the Hamiltonian, we assume that the matrix $\sigma$ is independent of time $t$. Hence, Condition $(C3)$ proposes the use of a diagonal matrix $\sigma$, which is simultaneously symmetric and antisymmetric.
  • (C3)
    The diffusion matrix $\sigma(\theta,I)\in C^{2}(\mathbb{R}^{n}\times\mathbb{R}^{n},\mathbb{R}^{n\times n})$ is diagonal, positive definite, and bounded for all $t\in[0,1]$. Additionally, it is Lipschitz continuous with respect to the second and third variables, with Lipschitz constant $L$. Furthermore, $\sigma$ satisfies a uniform ellipticity condition: there exist constants $0<\lambda\leq\Lambda$ such that
    $\lambda\leq\|\sigma\|_{HS}\leq\Lambda,\quad\forall t\in[0,1],$
    where $\|\cdot\|_{HS}$ denotes the Hilbert–Schmidt norm. Consequently, the inverse $\sigma^{-1}$ exists and is continuous, bounded, and symmetric for all $t\in[0,1]$.
@@
Assuming that Conditions $(C1)$ and $(C3)$ are satisfied, we further demonstrate, based on the established Onsager-Machlup functional, that the maximum probability path of the stochastic Hamiltonian system (2) satisfies the following nearly integrable Hamiltonian system:
$\dot{\varphi}(t)=J\nabla H(\varphi_{t})-\frac{1}{2}J\nabla H_{\sigma}(\varphi_{t}),$ (22)
where $H_{\sigma}(\varphi_{t}):=\frac{1}{2}\sum_{i=1}^{n}\left(\sigma_{\theta,i}(\theta,I)^{2}-\sigma_{I,i}(\theta,I)^{2}\right)$, $\sigma_{\theta}(\theta,I)=\left(\sigma_{\theta,i}(\theta,I)\right)_{i=1}^{n}$ and $\sigma_{I}(\theta,I)=\left(\sigma_{I,i}(\theta,I)\right)_{i=1}^{n}$.